# Efficient Inference for Large Reasoning Models: A Survey

### v1

@@ 2.1 Large Reasoning Model
SOLAR (Li et al., 2025) SFT Qwen2VL-7B-Instruct Math
@@
LightThinker (Zhang et al., 2025) SFT DeepSeek-R1-Distill-Qwen-7B, DeepSeek-R1-Distill-LLaMA-8B Language Understanding, Math, Scientific, Commonsense, Logic
@@
IBPO (Yu et al., 2025) RL LLaMA-3.1-8B Math
Implicit Latent CoT ICoT-KD (Deng et al., 2023) SFT GPT-2 Small/Medium Math
@@
SoftCoT (Xu et al., 2025c) SFT LLaMA-3.1-8B-Instruct, Qwen2.5-7B-Instruct Math, Commonsense, Symbolic Reasoning
@@ 2.2 Efficiency Challenge in LRM Inference
A key driver of LRMs’ remarkable reasoning capabilities is the scaling of inference-time compute, which enables complex reasoning through long CoTs (Chen et al., 2025; Guo et al., 2025; Jaech et al., 2024; Muennighoff et al., 2025; Liu et al., 2025a). Compared to standard short CoTs (Wei et al., 2022), which are often shallow, heuristic-driven, and less generalizable (Sprague et al., 2025), long CoTs empower LRMs to tackle complex tasks such as advanced mathematics (Xu et al., 2025a) and medical question answering (Huang et al., 2025). However, this shift has also introduced the phenomenon of overthinking, where LRMs consume excessive inference tokens and reasoning steps even for simple problems, yielding only marginal performance improvements (Ma et al., 2024; Chen et al., 2024; Wu et al., 2025c). In real-world applications such as software engineering agents, overthinking has been found to negatively correlate with issue resolution rates (Cuadron et al., 2025). Moreover, LRMs’ reliance on inference-time scaling exposes them to overthinking attacks, where adversarial actors inject benign yet computationally intensive decoy problems (e.g., Sudoku puzzles) into the context for retrieval-augmented question answering, triggering substantial computational overhead (Kumar et al., 2025).
@@ 3.1 Explicit Compact CoT
CoT Compression. Succinct CoT representations streamline inference while preserving solution quality. Constrained-CoT (Nayab et al., 2024) and CoD (Xu et al., 2025b) confine intermediate reasoning to essential steps, ensuring brevity without losing critical information. Beyond simple compression, Sketch-of-Thought (SoT) (Aytes et al., 2025) uses a smaller “router” model to prompt the main LLM to generate sketches of reasoning, offering a concise yet cognitively inspired overview. InftyThink (Yan et al., 2025) decomposes complex tasks into bounded-length segments, creating intermediate summaries at each step. Other frameworks adapt their compression mechanisms in real time. LightThinker (Zhang et al., 2025) introduces special tokens that trigger the model to dynamically compress its ongoing thought process, reducing redundancy. TALE-EP (Han et al., 2024) dynamically adjusts the allotted reasoning tokens depending on task complexity, while Meta-Reasoner (Sui et al., 2025) applies a contextual multi-armed bandit to optimize efficiency.
@@
Fine-Tuning on Compact Reasoning Chains. Fine-tuning on compact reasoning data enables LRMs to internalize efficient inference behaviors while keeping performance. C3oT (Kang et al., 2024) leverages an LLM to generate condensed versions of long CoTs, preserving essential structure before jointly training models on both full and compressed chains. Skip Steps (Liu et al., 2024c) curates expert-validated answers with condensed steps and fine-tunes LLMs to mimic these concise reasoning paths. SOLAR (Li et al., 2025) fine-tunes LLMs using datasets annotated for both correctness and the effectiveness of the underlying task-specific reasoning topology, encouraging minimal yet complete logic flows. VARR (Jang et al., 2024) identifies redundant sentences in CoTs by analyzing their contribution to answer correctness, then fine-tunes models on the distilled, non-redundant reasoning processes. TokenSkip (Xia et al., 2025) prunes reasoning chains token-by-token based on importance, followed by fine-tuning across various compression ratios to balance brevity and precision. From a parameter space perspective, TALE-EP (Han et al., 2024) enhances token-budget awareness via SFT and direct preference optimization (DPO). CoT-Valve (Ma et al., 2025) discovers a latent direction that controls reasoning length, enabling models to flexibly adjust their level of detail based on task demands.
@@
Reward-Based Incentivization. A growing body of work introduces explicit reward signals to reduce unnecessary CoT complexity while preserving accuracy. Kimi k1.5 (Kimi Team et al., 2025) integrates length-based rewards to discourage verbose reasoning. Similarly, O1-Pruner (Luo et al., 2025) detects “length disharmony” and applies harmonizing penalties that promote brevity without sacrificing solution quality. Arora et al. (Arora & Zanette, 2025) use reinforcement learning to train models that dynamically allocate computational resources based on task difficulty, balancing cost and precision. DAST (Shen et al., 2025b) proposes a Token Length Budget metric that aligns task complexity with output length, encouraging efficiency through targeted penalties and rewards. IBPO (Yu et al., 2025) adopts a constrained RL framework to control the distribution of reasoning across response groups based on inference cost. MRT (Qu et al., 2025) applies meta-reinforcement learning to balance exploration of novel reasoning paths with the exploitation of concise, proven ones. Recent approaches also explore interactive or user-directed mechanisms for length control. Claude 3.7 (Anthropic, 2025), the first hybrid reasoning model, introduces an extended thinking mode where users can prescribe token budgets. L1 (Aggarwal & Welleck, 2025) generalizes this idea with Length Controlled Policy Optimization (LCPO), enabling fully configurable CoT lengths at inference time.
@@ 3.2 Implicit Latent CoT
Heima(Shen et al., 2025a) condenses CoT stages into latent thinking tokens and incorporates an explanatory prompt at the decoder stage to interpret the compressed reasoning. SoftCoT (Xu et al., 2025c) utilizes a small instruction-tuned 1B model to obtain instance-specific latent thought tokens and trains a projection layer to incorporate thought tokens into LLM input. Token-Assorted CoT(Su et al., 2025) mixes latent and text tokens, encoding the initial part of the CoT into VAE-based discrete latent tokens while preserving the remainder as natural language, resulting in a hybrid representation that enhances reasoning efficiency.
@@
LightThinker (Zhang et al., 2025) zero-shot,tho. 90.14% DeepSeek-R1-Distill-Qwen-7B -
@@
SoftCoT (Xu et al., 2025c) zero-shot 85.81% Qwen2.5-7B-Instruct -
@@ 5.2 Trade-off Between Interpretability and Efficiency of Reasoning
To address this trade-off more effectively, future research should focus on developing adaptive inference strategies that optimize the balance between reasoning efficiency and interpretability. One promising direction is the integration of external verification mechanisms, such as symbolic reasoning (Besold et al., 2021; Gaur & Saunshi, 2023; Sui et al., 2024) or retrieval-based justifications (Gao et al., 2023), which can provide post-hoc explanations for implicit reasoning models. Besides, new empirical studies are needed to quantify how different efficiency techniques impact both model accuracy and human trust, guiding the development of LRMs that are both efficient and interpretable in real-world scenarios.
@@ 5.3 Ensuring Safety of Efficient Reasoning
Although the existing methods improve the token efficiency of the LRMs, they may destroy the alignment of LRMs, increasing the potential safety risks, e.g., jailbreaking attacks (Liu et al., 2024d; He et al., 2025a) and privacy leakage (Li et al., 2023).
@@
To address this problem, one promising direction is to add safety constraints during the training process, like data filtering for the SFT/DPO data and designing the safety-related reward in RL training. Besides, the failure of current monitors may be due to LRMs’ ability being stronger than LLM-based guard models. Thus, it is worth designing stronger reasoning-based safeguard models (Liu et al., 2025b) to monitor the training data or LRMs.
@@ 5.4 Broader Application of Efficient Reasoning
As shown in Table 1, existing LRMs are primarily applied in math (Xia et al., 2025; Li et al., 2025), code (Kimi Team et al., 2025), or AI research (OpenAI, 2025a) scenarios.
@@ 6.1 New Architecture
Hybrid Autoregressive and Diffusion Models. The fundamental limitation of autoregressive models is their sequential nature, which makes inference slow, particularly for reasoning tasks that require long chains of intermediate steps. A potential solution is integrating diffusion models into LRMs (Nie et al., 2025). Diffusion models generate entire sequences in parallel, allowing for global reasoning structure optimization rather than token-by-token generation. But the challenge lies in controlling the generated reasoning steps to ensure logical consistency. A promising direction is hybrid architectures that use autoregression for fine-grained control over reasoning while leveraging diffusion-based sampling for efficiency, enabling LRMs to reason in a structured yet accelerated manner.
@@
Memory-Efficient Transformer Variants. One of the primary inefficiencies in LRMs stems from the quadratic complexity of self-attention. Applying linear attention mechanisms(e.g., RWKV (Peng et al., 2023)) or state-space models (e.g., Mamba (Gu & Dao, 2023)) could drastically reduce memory consumption and improve inference speed. The challenge is that such architectures often struggle with long-range dependencies, which are crucial for reasoning. A key question is whether hybrid models can selectively apply full attention for critical reasoning steps while using approximate attention elsewhere to optimize efficiency.
@@
Graph-Based Reasoning Models. Autoregressive LRMs process information sequentially, generating one token at a time. While effective, this approach struggles with complex multi-step reasoning tasks where different pieces of information must be retrieved, combined, and reasoned over in a structured manner. Graph-based reasoning models (Besta et al., 2024) offer a structured alternative to autoregressive LRMs by representing reasoning as a graph, where nodes encode intermediate steps and edges define logical dependencies. This enables parallel exploration of multiple reasoning paths, reducing inefficiencies inherent in sequential token generation. Monte Carlo Tree Search (MCTS) (Silver et al., 2017) enhances this by adaptively expanding promising reasoning branches while pruning redundant ones, optimizing both token efficiency and inference time. However, key challenges include designing training objectives for graph construction, optimizing traversal heuristics, and balancing interpretability with efficiency. Future research should explore hybrid architectures that fuse explicit graph reasoning with neural representations (He et al., 2024b; 2025b; Zhou et al., 2023), ensuring both scalability and robust logical coherence.

### v2

@@ 2.1 Large Reasoning Model
TS (Zhang et al., 2025b) Intervention Qwen-2.5-7B/14B/32B Math
Fractured Sampling (Liao et al., 2025) Inference-time Scaling DeepSeek-R1/Qwen-1.5B/7B/14B Math, Scientific, Logic
RPC (Song et al., 2025) KV Cache Compression QwQ-32B/DeepSeek-R1-Distill-Qwen-7B Math, Code, Instruction
ThinkLess (Li et al., 2025b) Prompt Qwen-2.5-7B/14B, LLaMA3.1-8B Math, Commonsense, Logic, Scientific
PLAN-AND-BUDGET (Lin et al., 2025a) Prompt DeepSeek-R1-Distill-Qwen-32B, QwQ-32B, DeepSeek-R1-Distill-LLaMA-70B, OpenAI o4-mini Math, Instruction, Planning
TrimR (Lin et al., 2025b) Prompt Pangu Pro MoE, Pangu-R-38B, QwQ-32B, DeepSeek-R1-Distill-Qwen-32B Math, Scientific
SOLAR (Li et al., 2025a) SFT Qwen2VL-7B-Instruct Math
@@
LightThinker (Zhang et al., 2025a) SFT DeepSeek-R1-Distill-Qwen-7B, DeepSeek-R1-Distill-LLaMA-8B Language Understanding, Math, Scientific, Commonsense, Logic
@@
IBPO (Yu et al., 2025b) RL LLaMA-3.1-8B Math
LS-Mixture SFT (Yu et al., 2025a) SFT Qwen2.5-32B-Instruct Math
ConCISE (Qiao et al., 2025) SFT, SimPO DeepSeek-R1-Distill-Qwen-7B/1.5B Math, Reasoning
Elastic Reasoning (Xu et al., 2025e) RL E1-Math-1.5B/E1-Code-14B Math, Code
S-GRPO (Dai et al., 2025) RL Qwen3-8B/14B, DeepSeek-R1-Distill-Qwen-7B/14B Math, Scientific
TLDR (Zhang et al., 2025b) RL Qwen-2.5-7B/14B/32B Math
Adaptive GoGI-Skip (Zhuang et al., 2025) SFT Gemma3-1B/4B/12B, Qwen2.5-3B/7B Math
SelfBudgeter (Li et al., 2025c) RL DeepSeek-R1-Distill-Qwen-1.5B Math
Long Short (Ning et al., 2025) SFT, RL Qwen2.5-7B, Llama3.1-8B Math, Logical
Length-Aware Optimization (Yuan et al., 2025a) RL Qwen-2.5-7B Math, Logic
Prune-on-Logic (Zhao et al., 2025) SFT DeepSeek-R1-Distill-Llama-8B, DeepSeek-R1-Distill-Qwen7B Math, Logic
DRP (Jiang et al., 2025b) SFT DeepSeek-R1-Distill-Qwen-7B/1.5B Math
FlashThink (Jiang et al., 2025a) Prompt, SFT Qwen2.5, Llama-3.1-8B-Instruct, Mistral-7B-Instruct-v0.3, Qwen3 Math, Reasoning
AnytimeReasoner (Qi et al., 2025) RL DeepSeek-R1-Distill-Qwen-1.5B Math
VeriThinker (Chen et al., 2025b) SVFT DeepSeek-R1-Distill-Qwen-7B/14B, DeepSeek-R1-Distill-Llama-8B Math, Reasoning
LASER (Liu et al., 2025b) RL DeepSeek-R1-Distill-Qwen-1.5B/7B/32B Math, Reasoning, Code
R1-Compress (Wang et al., 2025b) SFT Qwen2.5-14B/32B-Instruct Math, Logic, Scientific
ACPO (Cheng et al., 2025) SFT, RL DeepSeek-R1-Distill-Qwen-1.5B/7B, DeepSeek-R1-Distill-Llama-8B Math
ConciseRL (Dumitru et al., 2025) RL DeepSeek-R1-Distill-Qwen-1.5B, STILL-3-1.5B-preview Math, Commonsense
CTS (Yuan et al., 2025b) STF Qwen2.5-7B/14B Math
PIR (Xiao et al., 2025) SFT Qwen-2.5-32B Math, Science
ConciseR (Song & Zheng, 2025) RL Qwen2.5-Math-7B Math
CoThink (Fan et al., 2025) SFT, RL, Distillation Qwen2.5-Instruct-32B, DAPO, DeepSeek-R1-Distill, QwQ Math
DTO (An et al., 2025) SimPO DeepSeek-R1-Distill-Qwen-1.5B, DeepScaleR-1.5B-Preview, Llama - 3.3 - 70B - Instruct Math, Reasoning
A*-Thought (Xu et al., 2025c) SFT QwQ-32B, DeepSeek-R1-Distill-Qwen-32B, s1.1-32B Math
Implicit Latent CoT Soft Thinking (Zhang et al., 2025c) Decoding Qwen-32B/70B, LLaMA-70B Math, Code
ICoT-KD (Deng et al., 2023) SFT GPT-2 Small/Medium Math
@@
SoftCoT (Xu et al., 2025d) SFT LLaMA-3.1-8B-Instruct, Qwen2.5-7B-Instruct Math, Commonsense,Reasoning
CoLaR (Tan et al., 2025) SFT, RL Llama-3.2-1B-Instruct/Qwen-1.5B Math
@@ 2.2 Efficiency Challenge in LRM Inference
A key driver of LRMs’ remarkable reasoning capabilities is the scaling of inference-time compute, which enables complex reasoning through long CoTs (Chen et al., 2025a; Guo et al., 2025; Jaech et al., 2024; Muennighoff et al., 2025; Liu et al., 2025a). Compared to standard short CoTs (Wei et al., 2022), which are often shallow, heuristic-driven, and less generalizable (Sprague et al., 2025), long CoTs empower LRMs to tackle complex tasks such as advanced mathematics (Xu et al., 2025a) and medical question answering (Huang et al., 2025). However, this shift has also introduced the phenomenon of overthinking, where LRMs consume excessive inference tokens and reasoning steps even for simple problems, yielding only marginal performance improvements (Ma et al., 2024; Chen et al., 2024; Wu et al., 2025c). In real-world applications such as software engineering agents, overthinking has been found to negatively correlate with issue resolution rates (Cuadron et al., 2025). Moreover, LRMs’ reliance on inference-time scaling exposes them to overthinking attacks, where adversarial actors inject benign yet computationally intensive decoy problems (e.g., Sudoku puzzles) into the context for retrieval-augmented question answering, triggering substantial computational overhead (Kumar et al., 2025).
@@ 3.1 Explicit Compact CoT
CoT Compression. Succinct CoT representations streamline inference while preserving solution quality. Constrained-CoT (Nayab et al., 2024) and CoD (Xu et al., 2025b) confine intermediate reasoning to essential steps, ensuring brevity without losing critical information. Beyond simple compression, Sketch-of-Thought (SoT) (Aytes et al., 2025) uses a smaller “router” model to prompt the main LLM to generate sketches of reasoning, offering a concise yet cognitively inspired overview. InftyThink (Yan et al., 2025) decomposes complex tasks into bounded-length segments, creating intermediate summaries at each step. Other frameworks adapt their compression mechanisms in real time. LightThinker (Zhang et al., 2025a) introduces special tokens that trigger the model to dynamically compress its ongoing thought process, reducing redundancy. TALE-EP (Han et al., 2024) dynamically adjusts the allotted reasoning tokens depending on task complexity, while Meta-Reasoner (Sui et al., 2025) applies a contextual multi-armed bandit to optimize efficiency.
@@
Fine-Tuning on Compact Reasoning Chains. Fine-tuning on compact reasoning data enables LRMs to internalize efficient inference behaviors while keeping performance. C3oT (Kang et al., 2024) leverages an LLM to generate condensed versions of long CoTs, preserving essential structure before jointly training models on both full and compressed chains. Skip Steps (Liu et al., 2024c) curates expert-validated answers with condensed steps and fine-tunes LLMs to mimic these concise reasoning paths. SOLAR (Li et al., 2025a) fine-tunes LLMs using datasets annotated for both correctness and the effectiveness of the underlying task-specific reasoning topology, encouraging minimal yet complete logic flows. VARR (Jang et al., 2024) identifies redundant sentences in CoTs by analyzing their contribution to answer correctness, then fine-tunes models on the distilled, non-redundant reasoning processes. TokenSkip (Xia et al., 2025) prunes reasoning chains token-by-token based on importance, followed by fine-tuning across various compression ratios to balance brevity and precision. From a parameter space perspective, TALE-EP (Han et al., 2024) enhances token-budget awareness via SFT and direct preference optimization (DPO). CoT-Valve (Ma et al., 2025) discovers a latent direction that controls reasoning length, enabling models to flexibly adjust their level of detail based on task demands.
@@
Reward-Based Incentivization. A growing body of work introduces explicit reward signals to reduce unnecessary CoT complexity while preserving accuracy. Kimi k1.5 (Kimi Team et al., 2025) integrates length-based rewards to discourage verbose reasoning. Similarly, O1-Pruner (Luo et al., 2025) detects “length disharmony” and applies harmonizing penalties that promote brevity without sacrificing solution quality. Arora et al. (Arora & Zanette, 2025) use reinforcement learning to train models that dynamically allocate computational resources based on task difficulty, balancing cost and precision. DAST (Shen et al., 2025b) proposes a Token Length Budget metric that aligns task complexity with output length, encouraging efficiency through targeted penalties and rewards. IBPO (Yu et al., 2025b) adopts a constrained RL framework to control the distribution of reasoning across response groups based on inference cost. MRT (Qu et al., 2025) applies meta-reinforcement learning to balance exploration of novel reasoning paths with the exploitation of concise, proven ones. Recent approaches also explore interactive or user-directed mechanisms for length control. Claude 3.7 (Anthropic, 2025), the first hybrid reasoning model, introduces an extended thinking mode where users can prescribe token budgets. L1 (Aggarwal & Welleck, 2025) generalizes this idea with Length Controlled Policy Optimization (LCPO), enabling fully configurable CoT lengths at inference time.
@@ 3.2 Implicit Latent CoT
Heima(Shen et al., 2025a) condenses CoT stages into latent thinking tokens and incorporates an explanatory prompt at the decoder stage to interpret the compressed reasoning. SoftCoT (Xu et al., 2025d) utilizes a small instruction-tuned 1B model to obtain instance-specific latent thought tokens and trains a projection layer to incorporate thought tokens into LLM input. Token-Assorted CoT(Su et al., 2025) mixes latent and text tokens, encoding the initial part of the CoT into VAE-based discrete latent tokens while preserving the remainder as natural language, resulting in a hybrid representation that enhances reasoning efficiency.
@@
LightThinker (Zhang et al., 2025a) zero-shot,tho. 90.14% DeepSeek-R1-Distill-Qwen-7B -
@@
SoftCoT (Xu et al., 2025d) zero-shot 85.81% Qwen2.5-7B-Instruct -
@@ 5.2 Trade-off Between Interpretability and Efficiency of Reasoning
To address this trade-off more effectively, future research should focus on developing adaptive inference strategies that optimize the balance between reasoning efficiency and interpretability. One promising direction is the integration of external verification mechanisms, such as symbolic reasoning (Besold et al., 2021; Gaur & Saunshi, 2023; Sui et al., 2024b; a) or retrieval-based justifications (Gao et al., 2023), which can provide post-hoc explanations for implicit reasoning models. Besides, new empirical studies are needed to quantify how different efficiency techniques impact both model accuracy and human trust, guiding the development of LRMs that are both efficient and interpretable in real-world scenarios.
@@ 5.3 Ensuring Safety of Efficient Reasoning
Although the existing methods improve the token efficiency of the LRMs, they may destroy the alignment of LRMs, increasing the potential safety risks, e.g., jailbreaking attacks (Liu et al., 2024d; He et al., 2025a) and privacy leakage (Li et al., 2023; Wang et al., 2025a).
@@
To address this problem, one promising direction is to add safety constraints during the training process, like data filtering for the SFT/DPO data and designing the safety-related reward in RL training. Besides, the failure of current monitors may be due to LRMs’ ability being stronger than LLM-based guard models. Thus, it is worth designing stronger reasoning-based safeguard models (Liu et al., 2025c) to monitor the training data or LRMs.
@@ 5.4 Broader Application of Efficient Reasoning
As shown in Table 1, existing LRMs are primarily applied in math (Xia et al., 2025; Li et al., 2025a; Gong et al., 2025), code (Kimi Team et al., 2025), or AI research (OpenAI, 2025a) scenarios.
@@ 6.1 New Architecture
Hybrid Autoregressive and Diffusion Models. The fundamental limitation of autoregressive models is their sequential nature, which makes inference slow, particularly for reasoning tasks that require long chains of intermediate steps. A potential solution is integrating diffusion models into LRMs (Nie et al., 2025). Diffusion models generate entire sequences in parallel, allowing for global reasoning structure optimization rather than token-by-token generation. But the challenge lies in controlling the generated reasoning steps to ensure logical consistency. A promising direction is hybrid architectures that use autoregression for fine-grained control over reasoning while leveraging diffusion-based sampling for efficiency, enabling LRMs to reason in a structured yet accelerated manner. While diffusion offers potential speedups, the overhead of managing this synchronicity and the potential need for iterative refinement to correct logical inconsistencies might offset some gains, especially when compared to optimized autoregressive approaches like speculative decoding, which also aim to accelerate generation without sacrificing as much direct control. The actual trade-off between generation speed, resource consumption (as diffusion models can be computationally intensive to train and sample from), and the quality of reasoning remains an open research question.
@@
Memory-Efficient Transformer Variants. One of the primary inefficiencies in LRMs stems from the quadratic complexity of self-attention. Applying linear attention mechanisms(e.g., RWKV (Peng et al., 2023)) or state-space models (e.g., Mamba (Gu & Dao, 2023)) could drastically reduce memory consumption and improve inference speed. The challenge is that such architectures often struggle with long-range dependencies, which are crucial for reasoning. A key question is whether hybrid models can selectively apply full attention for critical reasoning steps while using approximate attention elsewhere to optimize efficiency. The practicality of such hybrids depends on the effective identification of ”critical” steps and the seamless integration of different attention mechanisms without introducing excessive architectural complexity or training instability. Compared to methods like quantization or sparse attention, which aim to reduce the cost of full attention, linear attention and state-space models represent a more fundamental architectural shift. The actual benefit will depend on whether the reduced memory and computation translate to tangible improvements in reasoning quality per unit of resource, especially for tasks demanding high fidelity.
@@
Graph-Based Reasoning Models. Autoregressive LRMs process information sequentially, generating one token at a time. While effective, this approach struggles with complex multi-step reasoning tasks where different pieces of information must be retrieved, combined, and reasoned over in a structured manner. Graph-based reasoning models (Besta et al., 2024) offer a structured alternative to autoregressive LRMs by representing reasoning as a graph, where nodes encode intermediate steps and edges define logical dependencies. This enables parallel exploration of multiple reasoning paths, reducing inefficiencies inherent in sequential token generation. Monte Carlo Tree Search (MCTS) (Silver et al., 2017) enhances this by adaptively expanding promising reasoning branches while pruning redundant ones, optimizing both token efficiency and inference time. However, key challenges include designing training objectives for graph construction, optimizing traversal heuristics, and balancing interpretability with efficiency. Future research should explore hybrid architectures that fuse explicit graph reasoning with neural representations (He et al., 2024c; 2025b; Zhou et al., 2023; He et al., 2024b; Hou et al., 2023), ensuring both scalability and robust logical coherence.